OpenStudy (anonymous):

a 5.6m wire at 20C a 2.8m section has a diameter of .65 mm and the other portion has a 1.5mm diameter. the current is 5.5A if the resistivity is 1.7x10^8 ohm m at 20 C what is E in each section? what is the potential drop across the wire?

OpenStudy (anonymous):

I get 2.9 X 10 ^(-4) vm^-1 for the .65mm section and 5.3x10^(-5) vm^-1 for the 1.5mm section I get 9.61x10^(-4) V for the potential drop.

OpenStudy (anonymous):

I used resistivity = E/J and J = I/A which gives me \[\frac{ \rho i }{ A } = \left| E \right|\]

OpenStudy (anonymous):

E = phi/A :O what is that forumla? :O use Ohms law!

OpenStudy (anonymous):

how would you use ohms law V=IR to find the e field in the different sections of the wire?

OpenStudy (anonymous):

\[V= IR \ or \\ J = \sigma E\] ;-)

OpenStudy (anonymous):

so is it not ok to use \[\rho I /A \] which is really just \[J = \sigma E \]

OpenStudy (anonymous):

I only stuck with those because I was given rho

OpenStudy (anonymous):

ohhh.. it was that i thought you wrote rho into I .. . i thought it was phi.. was wondering what phi was.. lol xD

OpenStudy (anonymous):

ok so lemme check ur answers!

OpenStudy (anonymous):

oh lol ok

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

wow.. I am getting some absurdly high answer \[v = iR = i \frac{\rho l}{A} = 5.5 \times \frac {1.7x10^8 \times 2.8}{\pi \times ( \frac {0.65 \times 10^-3}{2})^2}\] that gives me a ridiculously high number!

OpenStudy (anonymous):

is it 10 power 8 or 10 power -8 ?? your resisitivity?

OpenStudy (anonymous):

10 ^ (-8)

OpenStudy (anonymous):

sorry about that I missed the - earlier ><

OpenStudy (anonymous):

ok.. so plug in.. those values.. i get 0.79 V

OpenStudy (anonymous):

@sseebeck hello? :P

OpenStudy (anonymous):

is that your result for voltage drop over the entire length of the wire?

OpenStudy (anonymous):

no only the first section! 2.8m long!

OpenStudy (anonymous):

Im trying to figure out why my numbers are so far off from yours

OpenStudy (anonymous):

oh ok I have 2.9 x 10 ^ ( -4 ) v/m

OpenStudy (anonymous):

that is still wrong.. maybe you plugged in wrong!

OpenStudy (anonymous):

its 0.65 mm.. don't forget to convert that to meters!

OpenStudy (anonymous):

hmm I have \[E = \rho i / A = ((1.7\times 10^{-8}) (5.5\times10^{-3}))/ (3.22\times10^{-7}\]

OpenStudy (anonymous):

why 5.5 X 10^-3? its 5.5 A.. not mili A xD

OpenStudy (anonymous):

ahh there I was supposed to be 5.5 m A

OpenStudy (anonymous):

* it

OpenStudy (anonymous):

i'm apparently getting tired

OpenStudy (anonymous):

you .. you missy.. you need spectacles :-/ then your answers are right you just need to convert the field to potential difference thats it!

OpenStudy (anonymous):

I just did Ea*La+Eb*Lb = Va-Vb = V

OpenStudy (anonymous):

and you are right I do need glasses! I was worried about this half of the problem as the professor wrote (this is a difficult problem) at the end which the way I did it it wasn't difficult at all just a tiny bit of arithmetic.

OpenStudy (anonymous):

exactly lol.. its so easy peezy lemon squeezy. you nailed it..!

OpenStudy (anonymous):

oh excellent thanks so much for the help!

OpenStudy (anonymous):

No problem!